OpenStudy (anonymous):

Determine the values of x where the tangent line is horizontal for the function: 1/(x2−16)(x−7) The value(s) of x where the tangent line to the graph of the function is horizontal is(are)_. I know that the derivative of the f(x) is f′(x)=−3x2+14x−16/(x−7)^2(x2−16)^2... but I'm confused on where to do next

OpenStudy (anonymous):

where to *go ><

OpenStudy (psymon):

Well, just deal with the numerator. Anything that would cause a 0 in the denominator would just be giving you a spot that is not differentiable. So just focus on the numerator and see if you can factor that to find your zeros and hence, the horizontal tangent lines.

OpenStudy (psymon):

There are convenient values, by the way : )

OpenStudy (anonymous):

-16/3 and -2/3?

OpenStudy (anonymous):

wait no that's not right..

OpenStudy (psymon):

Think you can correct the error?

OpenStudy (anonymous):

i'm going over my work but i'm not sure what i'm doing wrong.

OpenStudy (anonymous):

WAIT i think i got it lol

OpenStudy (psymon):

So what do ya got? : )

hero (hero):

Just a question @jbeloy17.... have you tried graphing this first? I know you have to do it algebraically as well, but try graphing it first to gain a visual perspective.

OpenStudy (anonymous):

i got 2 and -8/9 this time but that's not right..

OpenStudy (psymon):

\[-3x^{2}+14x-16 = -(3x^{2}-14x+16) = -(3x^{2}-6x-8x+16)\]Now factoring by grouping: \[-((3x^{2}-6x)+(-8x+16))\] \[-(3x(x-2)-8(x-2))\] \[-(3x-8)(x-2)\]

OpenStudy (anonymous):

@Hero i graphed it already but i need the exact values

OpenStudy (anonymous):

okay.. i kinda get it.

OpenStudy (psymon):

Is something still unclear, though?

OpenStudy (anonymous):

no i get it. i tried to do borrow and payback, or just minusing the 16 from the beginning. but yeah i get it.

OpenStudy (anonymous):

now i just set -(3x-8)(x-2)=0?

OpenStudy (psymon):

Each factor, yes. 3x-8 = 0 x-2 = 0

OpenStudy (anonymous):

2 and 8/3

OpenStudy (psymon):

Bingo

OpenStudy (anonymous):

okay!... but i put it into my homework online, and it says that it's not the right answer

OpenStudy (anonymous):

@psymon idk what i'm doing wrong!

OpenStudy (psymon):

Im trying to check it. I personally do not see an inssue, but I could easily be missing something.

OpenStudy (anonymous):

okay. thank you for your time!

OpenStudy (anonymous):

anything @psymon?

OpenStudy (psymon):

Not really. This is kinda random, but try \[\frac{ -7 \pm \sqrt{97} }{ 3 }\]

OpenStudy (psymon):

@jbeloy17 !!!!!You just made a sign mistake on the derivative x_x

OpenStudy (anonymous):

what?! ><

OpenStudy (anonymous):

.9496192673 and -5.616285934 ???

OpenStudy (psymon):

\[\frac{ 1 }{ (x^{2}-16)(x-7) }= (x^{2}-16)^{-1}(x-7)^{-1}\] \[-1(x^{2}-16)^{-2}(2x)(x-7)^{-1} + (x^{2}-16)^{-1}(-1)(x-7)^{-2}\] \[\frac{ -2x }{ (x^{2}-16)^{2}(x-7) }-\frac{ 1 }{ (x^{2}-16)(x-7)^{2} }\] \[\frac{ -2x(x-7)-(x^{2}-16) }{ (x^{2}-16)^{2}(x-7)^{2} }= \frac{ -2x^{2}+14x-x^{2}+16 }{ (x^{2}-16)^{2}(x-7)^{2} }\] \[\frac{ -3x^{2}+14x+16 }{ (x^{2}-16)^{2}(x-7)^{2} }\]\[\frac{ -14 \pm \sqrt{14^{2} - 4(-3)(16)} }{ -6 } = \frac{ -14 \pm 2\sqrt{97} }{ -6 }= \frac{ -7 \pm \sqrt{97} }{ 3 }\]

OpenStudy (psymon):

Leave it in radical form, dont turn it into decimals unless it instructsyou to.

OpenStudy (anonymous):

it says that i have to find the values

OpenStudy (psymon):

Well, those are the values. Theyre the EXACT values. Turning it into a decimal just turns the answer into an approximation.

OpenStudy (anonymous):

i put the answers in and it says that it's incorrect

OpenStudy (psymon):

Then try the decimals I suppose. But thats correct.

OpenStudy (psymon):

Unless you see an error in my math above :/

OpenStudy (anonymous):

thanks for everything @psymon really it means a lot!!

hero (hero):

Did inputting decimals work for you @jbeloy17 ?

OpenStudy (psymon):

Yeah, np. Did what I could, lol.

OpenStudy (anonymous):

no it didn't

OpenStudy (anonymous):

nope

hero (hero):

lol, okay

OpenStudy (psymon):

What was your suspicion, hero?

hero (hero):

Just trying something, that's all. Pay me no attention.

OpenStudy (psymon):

Lol, alrighty then.

OpenStudy (psymon):

Is that the exact wording of the question? @jbeloy17

OpenStudy (anonymous):

Determine the values of where the tangent line is horizontal for the function: jbeloy17 Determine the values of x where the tangent line is horizontal for the function: 1/(x^2−16)(x−7) The value(s) of where the tangent line to the graph of the function is horizontal is(are

OpenStudy (anonymous):

sorry copied it twice

OpenStudy (psymon):

Alright, just making sure in case there could have been a misinterpreting of the question.